OpenStudy (anonymous):

Help please....Use a Double- or Half-Angle Formula to solve the equation in the interval [0, 2π). (Enter your answers as a comma-separated list. Round your answers to three decimal places where appropriate.) sin 2θ + sin θ = 0

OpenStudy (anonymous):

start with \[2\sin(x)\cos(x)+\sin(x)=0\] then factor as \[\sin(x)(2\cos(x)+1)=0\] and finally set \[\sin(x)=0\] and solve for \(x\), also \[2\cos(x)+1=0\] and solve for \(x\)

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

yw